# Macroscopic entanglement between localized domain walls inside a cavity

### v1

@@ IV Macroscopic entanglement
While entanglement between the photonic and DW modes arise from the direct light-matter coupling in the linearized Hamiltonian in Eq. (5), the generation of entanglement between the two macroscopic DW modes in the absence of any nonlinearity is not straightforward. In our work, this is achieved by using strong coupling to move to an dispersive regime in terms of the effective detuning $\tilde{\Delta}$ from the cavity modes. Mathematically this entails that $\tilde{\Delta}\gg\omega_{1},\kappa_{a}$. In this limit, we adiabatically eliminate the cavity mode using the steady-state value of the $k^{th}$ root and the relation with the DW mode fluctuations $a={i\alpha\sum_{j}g_{j}(b_{j}+b^{\dagger}_{j})}/{(i\tilde{\Delta}^{(k)}-\kappa_{a})}$. The resulting Hamiltonian (derived in Appendix C) is given by
@@
Further qualitative behavior of entanglement and squeezing in the collective modes can be ascertained by the noise spectrum $S(\omega)$ of the optical mode as shown in Fig. 4. This can be experimentally detected [56] from the light reflected from the cavity as shown in Fig. 1. The calculation of the noise spectrum is shown in Appendix D. The spectrum exhibits level attraction with eigenvalues merging to form exceptional points (EPs) [57, 58] in one of the roots, while level repulsion in another, leading to the formation of entangled hybrid modes of photons and DWs. In Fig. 4(a) for root $k=1$ we find level attraction in the polaritonic modes formed with a cavity and a collective bright DW mode as they merge just before the resonance $\tilde{\Delta}=-\omega_{1}$ and separate in the large detuning regime. This leads to two second-order exceptional points. First, at the point where the two modes merge with square root dependency just before resonance $\tilde{\Delta}=-\omega_{1}$, and the second, where two modes separate again in the large detuning regime. The onset of these two EPs is also associated with the sudden discontinuity in the smallest eigenvalue $\nu^{\rm{min}}_{a|1(2)}$ and the associated two-mode squeezing, as shown in Fig. 3(c). There also exists a dark mode at $\omega=\omega_{1(2)}$, seen as a horizontal line at $\omega=\omega_{1}$ in Figs. 4(a)-(b), which is decoupled from the cavity [59], shows no spectral peaks, and remains invariant with the detuning.
A method to understand the origin of EPs is to consider an effective non-hermitian Hamiltonian. This is obtained by rewriting the linearized QLEs in the form of time-dependent Schrödinger-like equation for the linearized modes $i\partial_{t}u^{(k)}=\mathcal{H}u^{(k)}$, neglecting the input noise terms. The eigenvalues of $\mathcal{H}$ are directly related to the drift matrix eigenvalue set as $\{\lambda_{H}\}=i\{\lambda_{A}\}$. In particular, for the case of $\kappa_{a}=\kappa_{1(2)}$, we obtain exact analytical solutions for six eigenvalues of $\mathcal{H}$. Four of these represent collective bright modes
$\lambda^{(k)}_{\pm,j}=-i\kappa_{a}\pm\sqrt{\tilde{\Delta}^{(k)~2}_{\pm}+(-1)^{j}\Theta^{2}},$ (9)
where $\Theta^{2}=\sqrt{\tilde{\Delta}^{(k)~4}_{-}-8G^{2}\bar{n}_{a}^{(k)}\tilde{\Delta}^{(k)}\omega_{1}}$ and $\tilde{\Delta}^{(k)~2}_{\pm}=(\tilde{\Delta}^{(k)~2}\pm\kappa^{2}_{a})/2$.
@@
The other two eigenvalues $\lambda^{D}_{\pm}=-i\kappa_{1}\pm\omega_{1}$ are for non-interacting dark modes. The formation of second-order EPs occur when the term $\Theta$ inside the square root in Eq. (9) vanishes, leading to coalescence of eigenvalues and eigenvectors, at $\tilde{\Delta}^{(k)}=-\omega_{1}(2Gg_{1}\bar{x}_{a}^{(k)}/\kappa_{a})^{2}$ for $j=1$ branch. When $\kappa_{a}\neq\kappa_{1}$, the eigenvalues need to be numerically obtained from the drift matrix in Eq. (6), however, the dark mode eigenvalues retain the same value $\lambda^{D}_{\pm}$. While crossing these EPs, there are transitions from a region where the real part of two eigenvalues merge while the imaginary part splits as square root and vice versa. Such sudden transition of eigenvalues leads the squeezed state to jump from one stable root to another, leading to a jump in optimal squeezing we see in Fig. 3(c,d).
Similarly, in Fig. 4(b) for $k=2$, just close to the resonance $\tilde{\Delta}=-\omega_{1}$, the bright DW mode gets strong level repulsion and the cavity mode remains close to $\omega=0$ while the dark DW mode again remains invariant for all detunings. Thus, this resonance is associated with the maximum DW-photon entanglement $\mathcal{E}_{a|1(2)}$ as observed in Fig. 3(a). For $\omega_{1}\neq\omega_{2}$, we have two resonances at $\tilde{\Delta}=-\omega_{1}$ and $\omega_{2}=-10\omega_{1}$. For root $k=1$ as shown in Fig. 4(c), these two resonances thus result in two regimes of level attraction close to each of these resonances and are thus associated with four second-order EPs, each of which leads to discontinuity in the squeezing $\nu^{\rm{min}}_{a|1}$ and $\nu^{\rm{min}}_{a|2}$ as shown in Fig. 3(d). Similarly, for root $k=2$, we see the level repulsion at each of these resonances leading to maximum photon-DW entanglement $\mathcal{E}_{a|1(2)}$ at the respective resonances $\tilde{\Delta}=-\omega_{1(2)}$.
@@
An important aspect related to the robustness of the macroscopic entanglement between the two domain walls is thermal stability. In general, thermal fluctuations at finite temperatures are detrimental to entanglement in any dissipative system operating in the GHz regime. In quantum magnonics, these temperatures are typically in the milli-Kelvin regime [60, 61], which makes it experimentally challenging. Figure 5 shows the variation of entanglement between the two DWs with dissipation and temperature. Within a range of dissipation values, there exists a cutoff temperature $T_{c}$ beyond which thermal fluctuations prohibit any entanglement between the DWs. First, we note that altering the loss rates of either the photonic ($\kappa_{a}$) or DW modes and ($\kappa_{1}$) changes the amount of macroscopic entanglement but not the $T_{c}$. Remarkably, reducing $\kappa_{a}$ by a factor of 10 has the same positive effect on entanglement as that achieved by increasing $\kappa_{1}$ by a factor of 10. As such the entanglement between the DWs is dependent on the ratio $\kappa_{a}/\kappa_{1}$, as shown in Fig. 5(a), with the maximum achieved at 0.2. However, an improvement of the physically relevant cutoff temperature $T_{c}$ can be achieved by controlling the pinning potentials. This is due to the fact that number of thermal excitations in DW modes that can negate entanglement, is directly dependent on $\hbar\omega/k_{B}T$. In fact, an increase of the DW frequency $\omega_{1(2)}$ by an order of magnitude leads directly to an order of magnitude increase in the cut-off temperature $T_{c}$. As such, as observed in Fig. 5(b), within realizable experimental parameters, stable, robust entanglement can be sustained between the two domain walls up to 2 K, using DW frequencies of 100 GHz [62, 63].
@@ Acknowledgements.
We would like to thank K.G. Suresh for valuable discussions. R.G. acknowledges financial support from CSIR-HRDG, India in the form of SRF. H.S.D. acknowledges funding from SERB-DST, India under Core-Research Grant (CRG/2021/008918) and IRCC, IIT Bombay (RD/0521-IRCCSH0-001).
Appendix A Quantizing and capturing the inverse Faraday effect
@@
It was realized that when a DW oscillator is placed inside a cavity resonator, it can feel the radiation pressure forces from the electromagnetic field analogous to the optomechanical radiation pressure experienced by a suspended mirror in cavity optomechanics. The origin of such a force here is because of the magneto-optic coupling effect [67] between the electromagnetic field and the spin system. The permittivity tensor gets modulated by the local spin oscillations resulting in the interaction energy
@@
where $a_{R},~a_{L}$ are the cavity annihilation operators for the right- and left-circularly polarized ($\lambda=\pm 1$) states of light. The coupling strength is obtained as $g_{j}=fS^{j}_{\rm eff}\omega_{a}/4$. The dimensionless geometrical factor $S^{j}_{\rm eff}=x^{j}_{\rm zpf}A^{j}_{\perp}/V_{c}$ depends on the scattering cross-section $A^{j}_{\perp}$, the cavity volume $V_{c}$ and the zero point fluctuation $x^{j}_{\rm zpf}=\sqrt{\hbar/2m_{j}\omega_{j}}$ of the macroscopic DW, with mass $m_{j}=\rho V_{j}$, where $\rho$ is the density of the material and $V_{j}$ is the effective cross-sectional volume of the DW. We note that higher-order terms and the non-linear Kerr effect have been neglected. Experimentally, it is possible to use a chiral cavity resonator, where one of these polarizations can be fully suppressed [39], allowing us to deal with single-mode interaction $a=a_{R}$.
Appendix B Linearizing around steady states
@@
This will have either one or three solutions for mean photon number, we label them by $\bar{n}^{(k)}_{\alpha},~k\in\{1,2,3\}$. Now, we shift the Hamiltonian around these steady state values for each root by transformation $a\rightarrow\alpha+a,b_{j}\rightarrow\beta_{j}+b_{j}$ and we rewrite the $H^{(k)}$ in terms of these solutions similar to [51] and obtain:
$H^{(k)}_{2}$ $=-\tilde{\Delta}^{(k)}a^{{\dagger}}a+\sum_{j=1}^{N}\omega_{j}b_{j}^{{\dagger}}b_{j},$
$+\sum_{j=1}^{N}g_{j}\left(\alpha^{(k)}a^{{\dagger}}+\alpha^{*(k)}a\right)\left(b_{j}+b_{j}^{{\dagger}}\right),$
$H^{(k)}$ $=H^{(k)}_{2}+H_{3}.$ (26)
@@
Appendix C Adiabatic Elimination
@@
Substituting this back to the linear equation Eq. (27):
@@
Appendix D Output Optical Spectrum

### v2

@@ IV Macroscopic entanglement
While entanglement between the photonic and DW modes arise from the direct light-matter coupling in the linearized Hamiltonian in Eq. (5), the generation of entanglement between the two macroscopic DW modes in the absence of any nonlinearity is not straightforward. In our work, this is achieved by using strong coupling to move to an dispersive regime in terms of the effective detuning $\tilde{\Delta}$ from the cavity modes. Mathematically this entails that $\tilde{\Delta}\gg\omega_{1},\kappa_{a}$. In this limit, we adiabatically eliminate the cavity mode using the steady-state value of the $k^{th}$ root and the relation with the DW mode fluctuations $a={i\alpha\sum_{j}g_{j}(b_{j}+b^{\dagger}_{j})}/{(i\tilde{\Delta}^{(k)}-\kappa_{a})}$. The resulting Hamiltonian (derived in Appendix D) is given by
@@
Further qualitative behavior of entanglement and squeezing in the collective modes can be ascertained by the noise spectrum $S(\omega)$ of the optical mode as shown in Fig. 4. This can be experimentally detected [68] from the light reflected from the cavity as shown in Fig. 1. The calculation of the noise spectrum is shown in Appendix F. The spectrum exhibits level attraction with eigenvalues merging to form exceptional points (EPs) [69, 70] in one of the roots, while level repulsion in another, leading to the formation of entangled hybrid modes of photons and DWs. In Fig. 4(a) for root $k=1$ we find level attraction in the polaritonic modes formed with a cavity and a collective bright DW mode as they merge just before the resonance $\tilde{\Delta}=-\omega_{1}$ and separate in the large detuning regime. This leads to two second-order exceptional points. First, at the point where the two modes merge with square root dependency just before resonance $\tilde{\Delta}=-\omega_{1}$, and the second, where two modes separate again in the large detuning regime. The onset of these two EPs is also associated with the sudden discontinuity in the smallest eigenvalue $\nu^{\rm{min}}_{a|1(2)}$ and the associated two-mode squeezing, as shown in Fig. 3(c). There also exists a dark mode at $\omega=\omega_{1(2)}$, seen as a horizontal line at $\omega=\omega_{1}$ in Figs. 4(a)-(b), which is decoupled from the cavity [71], shows no spectral peaks, and remains invariant with the detuning.
@@
A method to understand the origin of EPs is to consider an effective non-hermitian Hamiltonian. This is obtained by rewriting the linearized QLEs in the form of time-dependent Schrödinger-like equation for the linearized modes $i\partial_{t}u^{(k)}=\mathcal{H}u^{(k)}$, neglecting the input noise terms. The eigenvalues of $\mathcal{H}$ are directly related to the drift matrix eigenvalue set as $\{\lambda_{\mathcal{H}}\}=i\{\lambda_{A}\}$. In particular, for the case of $\kappa_{a}=\kappa_{1(2)}$, we obtain exact analytical solutions for six eigenvalues of $\mathcal{H}$. Four of these represent collective bright modes
$\lambda^{(k)}_{\pm,j}=-i\kappa_{a}\pm\sqrt{\tilde{\Delta}^{(k)~2}_{{+}}+(-1)^{j}\Theta^{2}},$ (9)
where $\Theta^{2}=\sqrt{\tilde{\Delta}^{(k)~4}_{-}-8G^{2}\bar{n}_{a}^{(k)}\tilde{\Delta}^{(k)}\omega_{1}}$ and $\tilde{\Delta}^{(k)~2}_{\pm}=(\tilde{\Delta}^{(k)~2}\pm{\omega^{2}_{1}})/2$. The other two eigenvalues $\lambda^{D}_{\pm}=-i\kappa_{1}\pm\omega_{1}$ are for non-interacting dark modes. The formation of second-order EPs occur when the term $\Theta$ inside the square root in Eq. (9) vanishes, leading to coalescence of eigenvalues and eigenvectors, at $\tilde{\Delta}^{(k)}=-\omega_{1}(2Gg_{1}\bar{x}_{a}^{(k)}/\kappa_{a})^{2}$ for $j=1$ branch. When $\kappa_{a}\neq\kappa_{1}$, the eigenvalues need to be numerically obtained from the drift matrix in Eq. (6), however, the dark mode eigenvalues retain the same value $\lambda^{D}_{\pm}$. While crossing these EPs, there are transitions from a region where the real part of two eigenvalues merge while the imaginary part splits as square root and vice versa. Such sudden transition of eigenvalues leads the squeezed state to jump from one stable root to another, leading to a jump in optimal squeezing we see in Fig. 3(c,d).
Similarly, in Fig. 4(b) for $k=2$, just close to the resonance $\tilde{\Delta}=-\omega_{1}$, the bright DW mode gets strong level repulsion and the cavity mode remains close to $\omega=0$ while the dark DW mode again remains invariant for all detunings. Thus, this resonance is associated with the maximum DW-photon entanglement $\mathcal{E}_{a|1(2)}$ as observed in Fig. 3(a). For $\omega_{1}\neq\omega_{2}$, we have two resonances at $\tilde{\Delta}=-\omega_{1}$ and $\omega_{2}=-10\omega_{1}$. For root $k=1$ as shown in Fig. 4(c), these two resonances thus result in two regimes of level attraction close to each of these resonances and are thus associated with four second-order EPs, each of which leads to discontinuity in the squeezing $\nu^{\rm{min}}_{a|1}$ and $\nu^{\rm{min}}_{a|2}$ as shown in Fig. 3(d). Similarly, for root $k=2$, we see the level repulsion at each of these resonances leading to maximum photon-DW entanglement $\mathcal{E}_{a|1(2)}$ at the respective resonances $\tilde{\Delta}=-\omega_{1(2)}$. The analysis of the eigenvalues of the drift matrix $\{\lambda_{A}\}$ and EPs can also be extended to study correlations and entanglement at dissipative phases transitions. This is shown in Appendix E.
@@
An important aspect related to the robustness of the macroscopic entanglement between the two domain walls is thermal stability. In general, thermal fluctuations at finite temperatures are detrimental to entanglement in any dissipative system operating in the GHz regime. In quantum magnonics, these temperatures are typically in the milli-Kelvin regime [72, 73], which requires strong cooling efforts, making it experimentally challenging. Figure 5 shows the variation of entanglement between the two DWs with dissipation and temperature. Within a range of dissipation values, there exists a cutoff temperature $T_{c}$ beyond which thermal fluctuations prohibit any entanglement between the DWs. First, we note that altering the loss rates of either the photonic ($\kappa_{a}$) or DW modes and ($\kappa_{1}$) changes the amount of macroscopic entanglement but not the $T_{c}$. Remarkably, reducing $\kappa_{a}$ by a factor of 10 has the same positive effect on entanglement as that achieved by increasing $\kappa_{1}$ by a factor of 10. As such the entanglement between the DWs is dependent on the ratio $\kappa_{a}/\kappa_{1}$, as shown in Fig. 5(a), with the maximum achieved at 0.2. However, an increase of the physically relevant cutoff temperature $T_{c}$ can be achieved by controlling the pinning potentials. This is due to the fact that number of thermal excitations in DW modes that can negate entanglement, is directly dependent on $\hbar\omega/k_{B}T$. In fact, an increase of the DW frequency $\omega_{1(2)}$ by an order of magnitude leads directly to an order of magnitude increase in the cut-off temperature $T_{c}$. As such, as observed in Fig. 5(b), within realizable experimental parameters, stable, robust entanglement can be sustained between the two domain walls up to 2 K, using DW frequencies of 100 GHz [74, 75].
@@ Acknowledgements.
We thank K.G. Suresh for valuable discussions. R.G. acknowledges financial support from the Human Resource Development Group of CSIR, India, in the form of a Senior Research Fellowship. H.Y.Y. is supported by the National Key R& D Program of China (2022YFA1402700) and the National Natural Science Foundation of China (NSFC) (Grant No. 12574132). H.S.D. acknowledges funding from the Science and Engineering Research Board, Department of Science and Technology, India, under the Core Research Grant program (Grant No. CRG/2021/008918) and from the Industrial Research and Consultancy Centre, IIT Bombay (Grant No. RD/0521-IRCCSH0-001).
Data Availability
The codes and data that support the findings of this research article are openly available [^2] The source codes that support the findings of this article are openly available at: https://github.com/qid-iitb/Entanglement_DW, and include routines, and parameters required to reproduce the key results and figures..
Appendix A Quantizing the domain wall and capturing the magneto-optic coupling
@@
It was realized that when a DW oscillator is placed inside a cavity resonator, it can feel the radiation pressure forces from the electromagnetic field analogous to the optomechanical radiation pressure experienced by a suspended mirror in cavity optomechanics. The origin of such a force here is because of the magneto-optic coupling [80] between the electromagnetic field and the DW, arising from the inverse Faraday effect [59, 60]. The electric permittivity tensor gets modulated by the local spin oscillations resulting in the interaction energy
@@
where $a_{R},~a_{L}$ are the cavity annihilation operators for the right- and left-circularly polarized ($\lambda=\pm 1$) states of light. Both polarized modes exert a perpendicular force on the DW, but in opposite directions. However, the magnitude of the interaction remains the same and does not affect the system dynamics. Experimentally, it is possible to use a chiral cavity resonator, where one of these polarizations can be fully suppressed [47], allowing us to work with single-mode interaction. The coupling strength is obtained as $g_{j}=fS^{j}_{\rm eff}\omega_{a}/4$. The dimensionless geometrical factor $S^{j}_{\rm eff}=x^{j}_{\rm zpf}A^{j}_{\perp}/V_{c}$ depends on the scattering cross-section $A^{j}_{\perp}$, the cavity volume $V_{c}$ and the zero point fluctuation $x^{j}_{\rm zpf}=\sqrt{\hbar/2m_{j}\omega_{j}}$ of the macroscopic DW, with mass $m_{j}=\hbar^{2}/(K_{\perp}l\lambda_{\rm{DW}})$, where $l$ is the size of unit lattice and $K_{\perp}$ is the anisotropy constant of the material. We note that higher-order terms and the non-linear Kerr effect have been neglected.
We note that there are qualitative similarities and differences between Néel and Bloch DWs. The interaction between the DWs and the chiral cavity via the magneto-optic coupling is quite similar for both and can be modeled as in Sec. II. However, the direction of the radiation pressure is different for Néel walls, as the axis of rotation of spins is perpendicular to the plane of the DW. Importantly, higher DW oscillation frequencies can be obtained in Néel DWs [81]. However, the coupling is significantly weaker, and requires specialized geometry that allows for Dzyaloshinskii-Moriya interactions.
Appendix B System parameters
The physical parameters used to obtained the desired oscillation frequency of the Bloch domain walls and the magneto-optic coupling strength can be achieved in a physical implementation using iron garnets [82] with ytterbium (YIG) or bismuth and cerium substituted YIG variants [83, 84]), which are attractive for magneto-optic based interactions and entanglement, because they exhibit large Faraday rotation with very low magnetic damping. In fact materials such as Bi:YIG and Ce:YIG films show Faraday rotation that can be an order(s) of magnitude larger than plain YIG [85].
[…]
Appendix C Linearizing around steady states
@@
This will have either one or three solutions for mean photon number, we label them by $\bar{n}^{(k)}_{\alpha},~k\in\{1,2,3\}$. Now, we shift the Hamiltonian around these steady state values for each root [61], i.e., transformation $a\rightarrow\alpha+a,b_{j}\rightarrow\beta_{j}+b_{j}$, and rewrite $H^{(k)}$ (up to two operator terms) to obtain:
$H^{(k)}$ $=-\tilde{\Delta}^{(k)}a^{{\dagger}}a+\sum_{j=1}^{N}\omega_{j}b_{j}^{{\dagger}}b_{j},$
$+\sum_{j=1}^{N}g_{j}\left(\alpha^{(k)}a^{{\dagger}}+\alpha^{*(k)}a\right)\left(b_{j}+b_{j}^{{\dagger}}\right).$ (26)
@@
Appendix D Adiabatic Elimination
@@
Substituting Eq. (31) back to the linear equation Eq. (27):
@@
Figure 6: Entanglement near dissipative phase transition. (a) The plot shows the entanglement variation with effective coupling $G_{\text{eff}}/\omega_{1}$ and detuning $\tilde{\Delta}/\omega_{1}$, for a small region of Fig. 2(c). The blue dashed line is the bifurcation line. (b) The change in local and intermode covariance of photon and DW modes at $G_{\text{eff}}/\omega_{1}\approx 1.5$. (d) and (e) show the eigenvalues of the drift matrix $\{\lambda^{(k)}_{A}\}$ as it varies with effective coupling and detuning, respectively. These are changes across the green-dotted line in (a). The plots in (c) and (f) show the change in intermode covariance and entanglement, respectively, in the vicinity of the bifurcation line (green dashed line in Fig. 2(a)) as the detuning is $\tilde{\Delta}$ is increased. The plots highlight the critical detuning for onset of DW-DW entanglement, which is equal to the ratio of photon and DW mode loss rate $\kappa_{a}/\kappa_{1}$.
@@
Appendix E Correlations near the dissipative phase transition
To qualitatively understand the physical origin of entanglement across the bifurcation line, we investigate the eigenvalues of the drift matrix $\lambda_{A}$, similar to the analysis done in Sec. IV, and also the correlations between the different modes. While the former tells us about the nature of the phase transition that leads to divergence of mode covariances, the latter highlights the specific correlations that build up between the mode quadratures. Figure 6(a), shows the entanglement in a specific region of the phase space, close to the continuous dissipative phase transition line (blue-dashed line). The dissipative phase transition occurs when the steady state roots $k=0$ and $k=2$ exchange their stability. Figures 6(d,e) shows that the real part of $\lambda_{A}$ approaches zero as one gets close to the critical line by either changing the effective coupling $G_{\text{eff}}$ (Fig. 6(b)) or the dispersion $\tilde{\Delta}$ (in Fig. 6(c)). The square root shaped bifurcation gives rise to an exceptional point with a significant build up DW-DW entanglement. In fact, the entire transition line is an exceptional line [89], where a pair of eigenvalues coalesce to zero at each point, which results in large correlations. For instance, Fig. 6(b) shows the divergence of the intermode covariances $\langle x_{1}x_{2}\rangle$ and $\langle p_{a}x_{1}\rangle$, and the local mode fluctuations $\langle x_{1}x_{1}\rangle$ and $\langle p_{a}p_{a}\rangle$. The covariances show that the quadratures $x_{1}$ and $x_{2}$ are correlated and comparable to the local fluctuations of $x_{1}$ and $x_{2}$, which entangles the two DW modes. Moreover, the quadratures $p_{a}$ and $x_{1}$ are anticorrelated and comparable with $\langle p_{a}p_{a}\rangle$, leading to entanglement between photon and DW modes, as evident from calculations of logarithmic negativity.
Another important observation is the onset of DW-DW entanglement along the bifurcation line. Analyzing the covariances in Fig. 6(c), we see that the onset of DW-DW entanglement (Fig. 6(f)) corresponds to the point $\tilde{\Delta}/\omega_{1}=2$, where $\langle x_{1}x_{2}\rangle=\langle x_{1}x_{1}\rangle>\langle p_{a}x_{1}\rangle>\langle p_{a}p_{a}\rangle$. As such, the DW-DW entanglement increases as photo-DW entanglement decreases. Interestingly, Fig. 6(f)) shows that DW-DW entanglement along the transition line becomes non-zero at $\tilde{\Delta}/\omega_{1}=\kappa_{a}/\kappa_{1}$, where the effective dispersion $|\Delta/\omega_{1}|$ starts becoming larger than the ratio of photon and DW loss rates.
Appendix F Output optical spectrum